# Streaming Technologies and Serialization Protocols: Empirical Performance Analysis

### v1

@@ IV-B Dataset
In our experiments, we consider eight different payloads, ranging from simple data to common machine learning workloads, and include fusion science data. Our goal is to cover a range of scenarios. This section briefly describes the datasets used to evaluate performance with various streaming technologies and serialization protocols.
  • Numerical Primitives: As a baseline comparison, we use simple datasets consisting of randomly generated numerical primitives for int32, float32, and float64 types.
  • BatchMatrix: A synthetic dataset where each message consists of a randomly generated 3D tensor of type float32 with shape $\{32,100,100\}$ to simulate sending a batched set of image samples.
  • Iris Data: This is a dataset using the well-known Iris dataset [41]. The Iris dataset contains an array of four float32 features and a one-dimensional string target variable.
  • MNIST: We use the widely used MNIST machine learning image dataset [42] as a realistic example of streaming 2D tensor data.
  • Scientific Papers: The scientific papers dataset is a well-known dataset in the field of NLP and text processing [43]. The dataset comprises 349,128 articles of text from PubMed and arXiv publications. Each sample is repeated as a collection of string for properties such as article, abstract, and section names for transmission.
  • Plasma Current Data: A more realistic example of scientific data, we use plasma current data from the MAST tokamak [4]. Each set of plasma current data contains three 1D arrays of type float32: data, time, and errors. The “data” array represents the amount of current at each timestep, the “time” represents the time the measurement was taken in seconds, and the “errors” represents the standard deviation of the error in the measured current.
@@ IV-C Implementation and Experimental Setup
  • DataStream: component handles loading data for transmission. This can be any one of the payloads described in section IV-B.
  • Producer: functions as the server side of the application. It packages data from the selected data stream and transmits it over the wire using the selected streaming technology, which may be any of the technologies described in section III-B.
  • Marshaler: handles the serialization of the data from the stream using the specified serialization protocol. This can be any of those described in section III-A.
@@
  • Consumer: functions as the client side of the application. It receives data transmitted by the producer using the selected streaming technology, processes the incoming messages, and performs the necessary actions. Producers and consumers interact using a configured protocol.
  • Broker: required by the streaming protocol (e.g., for Kafka, RabbitMQ, etc.) are run externally from the test in the background. In our framework, we configure all brokers using docker-compose [47] to ensure that our broker configurations are reproducible for every test.
  • Logger: is used by the marshaler to capture performance metrics for each test in a JSON file. For each message sent, the logger captures four timestamps: 1) before serialization, 2) after serialization, 3) after transmission, and 4) after deserialization. Using these four timestamps, we can calculate the serialization, deserialization, transmission, and total time. Additionally, the logger captures the payload size of each message immediately after serialization. With this additional information, we can calculate the average payload size and throughput of the streaming service.
ADIOS and ZeroMQ can directly send array data without copying the input array. However, to achieve this, the array data must be directly passed to the communication library without serialization. Therefore, we additionally consider ZeroMQ and ADIOS to have their own ”native” encoding strategy for each stream, which is only used with their respective streaming protocol. This allows for a fair comparison with other technologies because sending an encoding array with ADIOS or ZeroMQ incurs an additional copy that could be circumvented by properly using their zero-copy functionality.
@@ V Results
In this section we present the results of our experiments with the combination of different streaming technologies, serialization protocols, and data streams.
1) Object Creation Latency – We use different datasets that originate from various data analysis types, like NumPy or Xarray dataset. Depending on the encoding protocol, we may need to copy the data from its native format to a specific format like Capn’Proto or Protobuf objects. This copying process adds some overhead that should be taken into consideration. However, for encoding protocols like JSON, BSON, and Pickle that do not require format changes, we store the data in a Pydantic class. The results in Figure 3 show that for larger array datasets like BatchMatrix, Plasma, and MNIST encoding methods such as Protobuf, Thrift, and Captn’Proto tend to have higher object creation latency as they need to copy data into their own data types.
@@
Figure 3: Object latency ($L_{o}$) measured in milliseconds, for various data streams (x-axis) and encoding methods (color).
2) Object Creation Throughput – We consider the object creation throughput for each serialization method. The object creation time measures the time to convert data from the native data structure (such as a NumPy array) to the serialization format. Object creation time is important to consider if the format that the data will be used in will be different from the format it is sent over the wire. Typically, object creation force a copy of the data to be sent, which impacts the total throughput, especially when considering large array like data.
@@
Figure 4 shows the object creation throughput for each dataset and each encoding method. It is interesting to note here that protocol based methods incur a greater penalty for the object creation. This effect is more noticeable in larger datasets such as the BatchMatrix and Plasma datasets.
@@
Figure 4: Object creation throughput ($T_{o}$) measured in megabytes per second, for various data streams (x-axis) and encoding methods (color).
@@
Figure 5: The compression ratio ($C$), for various data streams (x-axis) and serialization protocols (color).
@@
3) Compression Ratio – The payload size, a crucial performance metric of serialization protocols, is independent of the choice of streaming protocols. Therefore, we have calculated the average compression ratio over all runs for each serialization protocol. Figure 5 presents the results for each protocol and data stream. Notably, Pickle, Avro, and XML consistently produce the largest payload sizes, often exceeding the original size. This is due to the inefficiency of their text-based encodings and the additional meta-data tags they add as overhead. Pickle, a binary format for storing Python objects, is particularly known for its large sizes and is not optimal for encoding data for streaming.
@@
The results show that the serialization protocol, Capn‘Proto, outperforms others in terms of payload size. The packed option of Capn‘Proto, also known as capnp-packed, is responsible for additional size efficiency. The capnp-packed format is closely followed by several binary serialization formats that show similar performance. The reason behind this performance can be attributed to their ability to achieve near-identical compression, which is close to the limits of what is possible for that particular data stream.
@@
Examining across data streams, it can be seen that the BatchMatrix dataset is fundamentally limited. This is because it is made up of randomly generated numbers, making it incompressible due to the lack of redundancy in the data. Conversely, for more realistic data such as MNIST and Plasma, a much higher compression ratio is achieved.. Better compression is achieved for formats such as Capn’Proto Packed, which exploit the redundancy in the data to achieve greater compression. Text-based formats, such as YAML, JSON, XML, and Avro, achieve significantly worse compression in comparison. In fact, due to the extra markup required for these formats, they can produce a larger payload size that the original data.
@@
Figure 6: Serialization ($L_{s}$) latency for different data streams (x-axis) and encoding protocols (color). Protocol encoding methods such as Protobuf and Captn’Proto consistently offer the best performance in terms of both serialization. Text-based encodings (YAML, XML, etc.) add a large latency penalty to serialization by increasing the verbosity of the data.
@@
Figure 7: Deserialization ($L_{d}$) latency for different data streams (x-axis) and encoding protocols (color). As with serialization, protocol encoding methods such as Protobuf and Captn’Proto offer the best performance. Likewise, Text-based encodings (YAML, XML, etc.) add a large latency penalty to serialization by increasing the verbosity of the data.
@@
4) Serialization Latency – The results for serialization time are shown in figure 6. There is a clear trend across all data streams from text based protocols being the slowest (Avro, YAML, etc.) towards binary encoded protocol-based methods (Capn’Proto, protobuf, etc.) being the fastest. Binary encoded but no-protocol methods fall in between these two extremes.
@@
It is interesting to note that Capn’Proto has the fastest serialization time. This is likely due to the fact that Capn’Proto stores data in format that is ready for serialization over the wire.
@@
5) Deserialization Latency – The results for serialization time are shown in figure 7. Again, a clear trend may be seen across all data streams from text based protocols to binary encoded protocol based methods.
Like serialization, Capn’Proto is generally the fastest deserialization methods across all tests. As mentioned above, this is likely due to Capn’Proto storing the data in a pre-serialized form.
@@
6) Serialization Throughput – Figure 8 shows the average throughput for serialization of the data using different types of protocols. It is evident from the graph that serialization techniques based on protocols such as ProtoBuf, Thrift, and Capn’Proto offer the highest serialization throughput. Binary methods that are protocol independent offer moderate throughput performance with the added advantage of greater flexibility as compared to protocol methods. Text-based methods perform the worst due to their high serialization overhead.
@@
Surprisingly, Avro also performs quite well by this metric. We believe that despite being a human-readable text-based method, it is also a protocol-based method. This means that both the producer and consumer are aware of the types and data structures being transmitted over the wire, facilitating faster throughput.
@@
Figure 8: Serialization ($T_{s}$) and deserialization ($T_{d}$) throughput for each encoding averaged over all data streams.
@@
7) Deserialization Throughput – Figure 8 also shows the average throughput for deserialization of the data using different types of protocols. It is noticeable that deserialization throughput time across all methods is smaller, indicating that deserialization is a main bottleneck to transmission.
@@
8) Transmission Latency – Figure 9 shows the transmission latency for various combinations of serialization and streaming technologies. The heatmap in each combination is sorted by the average latency from lowest to highest for each streaming technology.
@@
Across all technologies, it is observed that transmission latency is largely dependent on the choice of streaming technology rather than the choice of serialization protocol. In streaming technologies, a broker is required as an intermediary, which increases the overall latency, whereas RPC technologies have no broker, and hence, have lower latency. Among messaging technologies, RabbitMQ performs better with larger payloads, while ActiveMQ achieves lower latency with smaller payloads but performs worst on the largest payload (e.g., BatchMatrix). In RPC-based methods, Thrift consistently has the lowest latency except for the BatchMatrix stream, where Capn’Proto narrowly beats Thrift.
@@
With larger payloads such as BatchMatrix and plasma data streams, the impact of serialization protocol becomes more noticeable. It is challenging to identify a trend between encoding protocols in terms of latency, except that it is crucial to note the inefficiency of using XML and YAML for larger payloads.
@@
For the BatchMatrix data stream, an issue arises when sending a large YAML-encoded payload through the Python API, which causes ADIOS to produce a segmentation fault. Therefore, subsequent latency and throughput plots result in NaN, the empty cells in Figure 9.
@@
Figure 9: Each subplot or heatmap shows transmission latency ($L_{trans}$) for different serialization protocols (vertical axis) and streaming technologies (horizontal axis). Dark red indicates higher latency.
@@
9) Transmission Throughput – By examining the throughput, we gain better understanding of how different protocols affect transmission. Figure 10 shows that RPC methods achieve higher transmission throughput than message streaming technologies. When dealing with larger payloads, such as the BatchMatrix and Plasma data streams, protocol-based serialization choices such as Thirft, Capn’Proto, and Protobuf provide higher throughput than other methods. Interestingly, messagepack also performs well with larger payloads.
@@
Similar to latency, the choice of streaming technology is more important than the encoding. However, a trend towards protocol encoding methods can be observed on some larger datasets, such as the plasma dataset.
@@
Figure 10: Each subplot or heatmap shows transmission throughput ($T_{trans}$) for different serialization protocols (vertical axis) and streaming technologies (horizontal axis).
@@
10) Total Latency – Figure 11 shows the total latency for all combinations tat were tested. As before, it is clear that the Thrift, Capn’Proto, and ZeroMQ all perform well in these tests. ZeroMQ offers the lower latency in the BatchMatrix test because it avoids the overhead of copying the data into a new structure, as is the case with Thift or Protobuf. Among the broker based methods, RabbitMQ consistently performs well.
@@
When it comes to encoding methods, protocol-based methods generally perform the best across all datasets and streaming methods. However, it is not clear which method offers the lowest latency in general. Protocol-based methods can achieve high throughput by mixing encoding protocols and RPC frameworks. For example, considering the MNIST dataset, Capn’Proto achieves the lowest latency with the thrift protocol.
@@
There is a clear trend towards protocol encoding for complex data sets such as Iris, MNIST, and Plasma. Among streaming technologies, Thrift generally shows the best performance.
@@
Figure 11: Each subplot or heatmap shows total latency ($L_{tot}$) for different serialization protocols (vertical axis), and streaming technologies (horizontal axis).
@@
11) Total Throughput – Figure 12 shows total throughput, which is consistent with the total latency results discussed in the previous section. Protocol-based methods achieve the highest throughput. Among all the serialization protocols, Thrift is generally the best performing one. ZeroMQ performs well with the biggest dataset, BatchMatrix. Although the best encoding method is inconclusive, it shows a trend toward protocol-based methods, which give the highest throughput.
@@
Figure 12: Each subplot or heatmap shows total throughput ($T_{tot}$) for different serialization protocols (vertical axis), and streaming technologies (horizontal axis).
@@
Figure 13: Total throughput $T_{tot}$ for various batch sizes (e.g., ranging from 1 to 256) for the MNIST dataset.
@@
12) Effect of Batch Size on Throughput – In machine learning applications, data is often processed in batches. Our findings underscore the potential of batching data before transmission to enhance throughput. However, it is crucial to note that the batch size can significantly impact the throughput of the method used.
Figure 13 shows the throughput of the MNIST dataset with a variable batch size. As the batch size increases, the throughput reduces. This is due to the increased overhead of copying and serializing data for transmission. However, when the batch size is increased beyond 32 images per batch, the overall throughput begins to improve because fewer packets are needed to be communicated over the network. For binary and protocol encoding methods, increasing the batch size is shown to also increase the throughput. This observation is consistent with previous results; generally, protocol-based methods offer the best throughput. At larger batch sizes ($>128$), the throughput continues to increase because fewer transmission time is significantly slower than the serialization/deserialization cost. So grouping many examples into a single transmission improves throughput.

### v2

@@ IV-B Dataset
In our experiments, we consider eight different payloads, ranging from simple data to common machine learning workloads, and includes data from the fusion science domain. Our goal is to cover a range of scenarios. This section briefly describes the datasets used to evaluate performance with various streaming technologies and serialization protocols.
  1. 1–3)
    Numerical Primitives refer to basic data types that represent fundamental numerical values, e.g., NumPy data types[44]. As a baseline comparison, we use randomly generated numerical primitives for (1) int32, (2) float32, and (3) float64.
  2. 4.
    BatchMatrix is a synthetic dataset where each message consists of a randomly generated 3D tensor of type float32 with shape $\{32,100,100\}$ to simulate sending a batched set of image samples. The random nature of the synthetic data makes it incompressible.
  3. 5.
    Iris Data is a dataset using the well-known Iris dataset [45]. The Iris dataset contains an array of four float32 features and a one-dimensional string target variable.
  4. 6.
    MNIST is the widely used machine learning image dataset [46] as a realistic example of streaming 2D tensor data.
  5. 7.
    Scientific Papers dataset is a well-known dataset in the field of NLP and text processing [47]. The dataset comprises 349,128 articles of text from PubMed and arXiv publications. Each sample is repeated as a collection of string for properties such as article, abstract, and section names for transmission.
  6. 8.
    Plasma Current Data is a realistic example of scientific data, we use plasma current data from the MAST tokamak [4]. Each item of plasma current data contains three 1D arrays of type float32: data, time, and errors. The data array represents the amount of current at each timestep, the time represents the time the measurement was taken in seconds, and the errorsrepresent the standard deviation of the error in the measured current.
@@ IV-C Implementation and Experimental Setup
  • DataStream component handles loading data for transmission. This can be any one of the payloads described in section IV-B.
  • Producer functions as the server side of the application. It packages data from the selected data stream and transmits it over the wire using the selected streaming technology, which may be any of the technologies described in section III-B.
  • Marshaler handles the serialization of the data from the stream using the specified serialization protocol. This can be any of those described in section III-A.
@@
  • Consumer functions as the client side of the application. It receives data transmitted by the producer using the selected streaming technology, processes the incoming messages, and performs the necessary actions. Producers and consumers interact using a configured protocol.
  • Broker required by the streaming protocol (e.g., for Kafka, RabbitMQ, etc.) are run externally from the test in the background. In our framework, we configure all brokers using docker-compose [51] to ensure that our broker configurations are reproducible for every test.
  • Logger is used by the marshaler to capture performance metrics for each test in a JSON file. For each message sent, the logger captures four timestamps: 1) before serialization, 2) after serialization, 3) after transmission, and 4) after deserialization. Using these four timestamps, we can calculate the serialization, deserialization, transmission, and total time. Additionally, the logger captures the payload size of each message immediately after serialization. With this additional information, we can calculate the average payload size and throughput of the streaming service.
ADIOS and ZeroMQ can directly send array data without copying the input array. However, to achieve this, the array data must be directly passed to the communication library without serialization. Therefore, we additionally consider ZeroMQ and ADIOS to have their own “native” serializtion strategy for each stream, which is only used with their respective streaming protocol. This allows for a fair comparison with other technologies because sending a serialized array with ADIOS or ZeroMQ incurs an additional copy that could be circumvented by properly using their native zero-copy functionality.
@@ V Results
In this section we present the results of our experiments with the combination of different streaming technologies, serialization protocols, and data types. In this study, we used datasets derived from various data analysis frameworks, e.g., NumPy, Xarray, etc.
1) Object Creation Latency –
Some serialization protocols, such as Cap’n Proto and Protocol Buffers (Protobuf), require data to be converted from its native format. This data conversion introduces a performance overhead. Conversely, serialization protocols such as JSON, BSON, and Pickle, which do not require format alterations, allow for direct storage of data within a Pydantic class structure. The former approach minimizes data manipulation and potentially reduces processing time.
Figure 3 shows object creation latency across different serialization methods. The results demonstrate that protocols such as Protobuf, Thrift, and Cap’n Proto exhibit higher object creation latencies for larger array datasets like BatchMatrix, plasma, and MNIST. This increased latency is attributed to the necessary data copying process, where data must be transferred into the protocol-specific data types.
@@
Figure 3: Object creation latency ($L_{o}$), measured in milliseconds (ms), of various data types arranged in the x-axis and serialization methods shown in colored bars
2) Object Creation Throughput –
Object creation throughput measures the time required to convert data from its native structure (such as a NumPy array) into a serialization format, normalized by the size of the data. This metric is important when the format used for processing the data differs from the format used for transmission. In such cases, object creation often necessitates copying the data, which can significantly affect overall throughput, particularly for large datasets.
@@
Figure 4 shows the object creation throughput for each dataset and each serialization method. Protocol based methods incur a higher penalty for the object creation. This penalty is more noticeable in larger datasets such as the BatchMatrix plasma datasets, and scientific papers.
@@
Figure 4: Object creation throughput, ($T_{o}$) measured in megabytes per second (MB/s), of various data types arranged in the x-axis and serialization methods shown in colored bars
@@
Figure 5: The compression ratio ($C$) of various data types arranged in the x-axis and serialization methods shown in colored bars
3) Compression Ratio – This performance metric quantifies the efficiency of serialization by measuring the ratio between the original data size and the size of the serialized payload. This metric remains unaffected by the choice of streaming protocols.
@@
Figure 5 shows the compression ratio of serialization protocols. Pickle, Avro, and XML consistently produce largest serialized payloads, often exceeding the original data size. This inefficiency is due to their text-based serialization and the additional metadata tags that contribute to overhead. Pickle, despite being a binary format designed for storing Python objects, is particularly inefficient in terms of size, making it suboptimal for data streaming.
@@
The result show that binary compression algorithms perform best amongst all options. capnp-packed, Protobuf, CBOR, BSON, UBJSON, Thrift are all competitive in terms of compression ratio. The reason behind this performance can be attributed to their ability to achieve near-identical compression, which is close to the limits of what is possible for that particular data stream.
@@
Examining across data types, we can see that the BatchMatrix dataset is fundamentally limited. This is because it is made up of randomly generated numbers, making it incompressible due to the lack of redundancy in the data. For more realistic data such as MNIST and plasma, more variety in compression ratio is achieved. Data redundancy can be exploited to achieve a better compression ratio. For example capnp-packed. Text-based formats, such as YAML, JSON, XML, and Avro, achieve significantly worse compression ratio. In fact, due to the extra markup required for these formats, they can produce a larger payload size that the original data.
@@
Figure 6: Serialization ($L_{s}$) latency of various data types arranged in the x-axis and serialization methods shown in colored bars. Protocol serialization methods such as Protobuf and Captn’Proto consistently offer the best performance in terms of both serialization. Text-based serializers (YAML, XML, etc.) add a large latency penalty to serialization by increasing the verbosity of the data.
@@
Figure 7: Deserialization ($L_{d}$) latency of various data types arranged in the x-axis and serialization methods shown in colored bars. As with serialization, protocol methods such as Protobuf and Captn’Proto offer the best performance. Likewise, Text-based serialization methods (YAML, XML, etc.) add a large latency penalty by increasing the verbosity of the data.
4) Serialization Latency –
@@
The results for serialization time are shown in Figure 6. We observed a clear trend across all data types, with text-based protocols (e.g., Avro, YAML, etc.) showing the slowest serialization time, while binary-encoded protocol-based methods (e.g., Capn’Proto, Protobuf, etc.) demonstrate the fastest. Binary-encoded methods without protocol support methods fall between these two extremes.
@@
Capn’Proto achieves the fastest serialization times among all methods. This performance advantage is likely due to Capn’Proto design, which stores data in a format that is ready for serialization.
5) Deserialization Latency –
@@
Figure 7 shows the results for deserialization latency. We see a clear trend across all data types, with text-based protocols displaying longer deserialization times compared to binary-encoded, protocol-based method.
Like serialization, Capn’Proto consistently achieves the fastest deserialization times across all tests. This superior performance is likely due to its design, which stores data in a format that is already optimized for serialization and deserialization, reducing the need for additional processing.
@@
6) Serialization Throughput – Figure 8 shows the average throughput for serialization of the data using various protocols. The results show that protocols-based (e.g., ProtoBuf, Thrift, and Capn’Proto) serialization techniques achieve the highest throughput. Binary methods that are protocol-independent achieve moderate throughput. Text-based methods perform the worst due to their high serialization overhead.
@@
Surprisingly, Avro also performs well, although it is human-readable text-based method. Its protocol-based nature likely contributes to this efficiency - this implies that both the producer and consumer are aware of the types and structures being transmitted, facilitating faster throughput.
@@
Figure 8: Serialization ($T_{s}$) and deserialization ($T_{d}$) throughput for each serializer averaged over all data types.
@@
7) Deserialization Throughput – Figure 8 also shows the average throughput for deserialization of the data using various protocols. The results show that deserialization throughput is consistently lower across all methods, suggesting that deserialization is a significant bottleneck in data transmission process.
@@
8) Transmission Latency – Figure 9 shows the transmission latency for various combinations of serialization and streaming technologies. In the heatmap, combinations are sorted by the average latency from lowest to highest for each streaming technology.
@@
The results show that the transmission latency mainly depends on the choice of streaming technology rather than serialization protocol.Streaming technologies, which require a broker as an intermediary, introduce latency. In contrast, RPC-based technologies, which operate without a broker, achieve lower latency. Among messaging technologies, RabbitMQ performs better with larger payloads, while ActiveMQ achieves lower latency for smaller payloads but struggles with largest payloads (e.g., BatchMatrix). In RPC-based methods, Thrift consistently delivers the lowest latency except for the BatchMatrix stream, where Capn’Proto narrowly beats Thrift.
@@
For larger payloads such as BatchMatrix and Plasma data types, the choice of serialization protocol becomes more noticeable. While it is difficult to identify a trend in latency across serialization protocols, it is clear that XML and YAML are inefficient for handling larger payloads.
@@
For the BatchMatrix data, an issue arises when attempting to send a large YAML-encoded payload through the Python API, which causes a segmentation fault in ADIOS. As a result, subsequent latency and throughput measurements result in NaN values, which are represented as empty cells in Figure 9.
@@
Figure 9: Each heatmap shows transmission latency ($L_{trans}$) for each combination of serialization protocols (vertical axis) and streaming technologies (horizontal axis). Dark red indicates higher latency. The left-to-right trend (changes in color of the heatmap) indicates that the choice of streaming technology has the most significant impact on latency.
@@
9) Transmission Throughput – Figure 10 shows that RPC methods achieve higher transmission throughput. When handling larger payloads, such as the BatchMatrix and plasma data, protocol-based serialization methods such as Thirft, Capn’Proto, and Protobuf deliver higher throughput. Interestingly, MessagePack also performs well with larger payloads.
@@
Similar to latency, the choice of streaming technology plays an important role than the choice of serialization method. However, a trend favouring protocol-based serialization methods emerges with some larger datasets, such as the plasma dataset.
@@
Figure 10: Each heatmap shows transmission throughput ($T_{trans}$) for different serialization protocols (vertical axis) and streaming technologies (horizontal axis). The left-to-right trend (changes in color of the heatmap) indicates that the choice of streaming technology has the most impact than serialization protocol.
@@
10) Total Latency – Figure 11 shows the total latency across various methods. As observed earlier, Thrift, Capn’Proto, and ZeroMQ perform well in this metric. ZeroMQ achieves the lowest latency in the BatchMatrix data as it avoids the overhead associated with copying data into a new structure, but is necessary with Thift or Protobuf. Among the broker-based methods, RabbitMQ consistently performs well.
@@
In terms of serialization methods, protocol-based methods generally perform the best across all datasets and streaming technologies. However, it is unclear which method offers the lowest latency overall. Protocol-based methods can achieve high throughput by integrating different serialization protocols with RPC frameworks. For example, in case of the MNIST dataset, Capn’Proto achieves the lowest latency when used with the thrift protocol.
@@
A clear trend emerges in favor of protocol serialization for complex datasets such as Iris, MNIST, and plasma. Among streaming technologies, Thrift generally shows the best performance.
@@
Figure 11: Each heatmap shows total latency ($L_{tot}$) for different serialization protocols (vertical axis), and streaming technologies (horizontal axis). The color variations in the heatmap, observed from top to bottom and left to right, indicate that both streaming technology and serialization protocol affect overall system performance; however, streaming technology has the most significant influence.
@@
11) Total Throughput – Figure 12 shows total throughput, which is consistent with the total latency results discussed earlier. Protocol-based methods achieve the highest throughput, with Thrift emerging as the best-performing serialization protocol. ZeroMQ performs particularly well with the large dataset, e.g., BatchMatrix and plasma. While no single serialization protocol conclusively outperforms the others, there is a trend favoring protocol-based serialization protocols, which consistently deliver the highest throughput.
The choice of streaming technology has a more significant impact that the choice of serialization on throughput. This is largely due to the difference between broker based, which have message delivery guarantees, and RPC based systems which do not.
@@
Figure 12: Each heatmap shows total throughput ($T_{tot}$) for different serialization protocols (vertical axis) and streaming technologies (horizontal axis). While both factors influence total throughput, streaming technology has a more significant impact.
@@
Figure 13: Total throughput ($T_{tot}$) for various batch sizes (ranging from 1 to 256) for the MNIST dataset.
@@
12) Effect of Batch Size on Throughput – In machine learning applications, data is often processed in batches.
We selected the MNIST dataset as an example for this test. Figure 13 shows the throughput of the MNIST dataset with a variable batch size.
In most cases, when the batch size is increased beyond 32 images per batch, the overall throughput begins to improve because fewer packets are needed to be communicated over the network.
At larger batch sizes ($>128$), the throughput continues to increase because fewer transmission time is significantly slower than the serialization/deserialization cost. So grouping many examples into a single transmission improves throughput.
Text-based protocols experience reduced throughput with increasing batch size due to their verbose nature, increased parsing complexity, and higher input-output demands. Conversely, binary and protocol-encoded formats benefit from increased batch size because they are optimized for compactness, machine efficiency, and effective use of network bandwidth. This contrast arises from the fundamental design trade-offs between human readability and machine efficiency. This observation is consistent with previous results; generally, protocol-based methods offer the best throughput.